OpenStudy (anonymous):

What is the sum of an 8-term geometric series if the first term is 13 and the last term is -3,639,168?

OpenStudy (callisto):

common ratio of terms \[= \sqrt[7]{-3,639,168/13} = -6\] Sum of GS \[=\frac{a(r^n-1)}{r-1} =\frac{13((-6)^8-1)}{(-6)-1} = ? \]